OpenStudy (anonymous):

What is the period of vibration for a wave with a frequency of 150 hertz

OpenStudy (asadkarim7):

frequency =1/period so period = 1/frequency 1/150 = 6.667x10 ^-3

OpenStudy (anonymous):

What is the period of vibration for a wave with a frequency of 0.50 hertz

OpenStudy (anonymous):

The period of a wave is the time for a particle on a medium to make one complete vibrational cycle. Period, being a time, is measured in units of time such as seconds, hours, days or years. The formula for period = 1/frequency.

OpenStudy (anonymous):

Basically, we substitute the values in to solve for period (T): T = 1/f T = 1/(50) T = 0.02 s